# DocR1: Evidence Page-Guided GRPO for Multi-Page Document Understanding

### v2

@@ Introduction
Documents are ubiquitous in everyday life, encompassing scanned forms, tables, charts, and PDFs. Consequently, document understanding is a critical capability for multimodal large language models (MLLMs). While recent MLLMs have achieved strong performance on single-page document tasks (Hu et al. 2024a; Zhou et al. 2024; Liu et al. 2024; Wang et al. 2023; Zhao et al. 2024; Feng et al. 2024), these tasks often cover a subset of real-world applications. In practice, many tasks involve multi-page documents—such as scientific papers, contracts, and reports—which require not only fine-grained visual understanding but also the ability to retrieve relevant pages and reason across dispersed content.
@@
Recent advances, including OpenAI’s o1 (Jaech et al. 2024) and DeepSeek-R1 (Guo et al. 2025), highlight the growing importance of advanced reasoning in LLMs. Parallel work has begun to explore reinforcement learning (RL) in MLLMs. For example, VisualRFT (Liu et al. 2025) introduced verifiable reward functions for perception tasks and applied GRPO (Shao et al. 2024) to enhance object detection and visual classification. Similarly, Vision-R1 (Huang et al. 2025) integrated GRPO with a PTST training strategy to improve mathematical reasoning capabilities. However, these methods are restricted to single-image inputs and focus on either perception or symbolic reasoning. For multi-image tasks, methods like Video-R1 (Feng et al. 2025) and VideoChat-R1 (Li et al. 2025) have applied GRPO in the video domain, incorporating customized reward functions to capture spatiotemporal dependencies. Nevertheless, applying RL to multi-page document understanding—where effective evidence retrieval and multi-hop reasoning are both critical—remains largely unexplored.
@@
MLLMs have recently advanced document understanding by enabling unified modeling of textual and visual information, without relying on traditional OCR engines. Several recent works have explored MLLM-based document understanding. UReader (Ye et al. 2023b) pioneered OCR-free, end-to-end instruction tuning across multiple tasks. The mPLUG-DocOwl series (Ye et al. 2023a; Hu et al. 2024a, b) improved structure awareness and computational efficiency through structural modeling and visual token compression. TextMonkey (Liu et al. 2024) introduced shifted window attention to preserve semantic continuity across image patches. DOGE (Zhou et al. 2024) proposed a document-oriented grounding framework with high-quality training data, while Doc-CoB (Mo et al. 2025) enabled dynamic region selection for focused, step-by-step reasoning. However, most methods are still limited to single-page documents. Our work targets more complex multi-page document tasks in real scenarios.
@@
RL has shown promise in enhancing the reasoning capabilities of LLMs. While early approaches relied on pretrained reward models, recent work such as DeepSeek-R1 (Guo et al. 2025) demonstrates that simple, rule-based rewards can provide scalable and verifiable supervision. Building on this insight, researchers have extended RL to multimodal large language models (MLLMs). VisualRFT (Liu et al. 2025) introduced visually grounded, verifiable reward functions for fine-grained classification and few-shot detection using GRPO. Vision-R1 (Huang et al. 2025) further leveraged GRPO with a PTST strategy to improve mathematical reasoning on a curated multimodal dataset. In the video domain, Video-R1 (Feng et al. 2025) proposed T-GRPO to explicitly incorporate temporal cues during training, marking the first systematic application of GRPO to video-based MLLMs. VideoChat-R1 (Li et al. 2025) also applied GRPO to enhance spatiotemporal reasoning, achieving strong performance on temporal grounding and tracking tasks. While these studies focus on perception and video reasoning, RL has not yet been explored for multi-page document understanding. Our work aims to fill this gap.
@@
While the GRPO algorithm (Shao et al. 2024) has demonstrated promise in enhancing reasoning capabilities, its direct application to multi-page document understanding remains suboptimal. To address this gap, we introduce Evidence Page-Guided GRPO (EviGRPO), a variant of GRPO tailored specifically for multi-page document tasks, as illustrated in Figure 2.
@@
As illustrated in Figure 3, we develop a rigorous two-stage data annotation pipeline. In the first stage, given an input and a task-specific prompt, we employ the Gemini 2.5 Flash model (Comanici et al. 2025) to generate a target output. The sample advances to the next stage only if its predicted answer is consistent with the ground truth. In the second stage, the same MLLM is prompted again with the annotated content in order to verify its accuracy. The annotation is retained only if the model’s output once again aligns with the ground truth under this controlled setting.
@@
As shown in Table 2, based on the proposed annotation pipeline, we presents the statistics of our annotated dataset. The dataset includes both single-page and multi-page document samples. Specifically, for single-page documents, we curate 13 widely-used datasets, including DocVQA (Mathew, Karatzas, and Jawahar 2021), InfoGraphicsVQA (Mathew et al. 2022), ChartQA (Masry et al. 2022), DeepForm (Svetlichnaya 2020), DVQA (Kafle et al. 2018), FigureQA (Kahou et al. 2017), KleisterCharity (Stanisławek et al. 2021), OCRVQA (Mishra et al. 2019), TabFact (Chen et al. 2019), TextCaps (Sidorov et al. 2020), TextVQA (Singh et al. 2019), VisualMRC (Tanaka, Nishida, and Yoshida 2021), and WikiTableQuestions (Pasupat and Liang 2015). These datasets cover a diverse range of document types. For instance, DocVQA and DVQA contain scanned pages and visual charts, while DeepForm, TabFact, and WikiTableQuestions focus on structured tabular formats. This diversity enables the model to learn from various layouts and visual structures commonly encountered in real-world documents. For the multi-page document domain, we integrate three widely used datasets: DUDE (Van Landeghem et al. 2023), MP-DocVQA (Tito, Karatzas, and Valveny 2023), and TATDoc (Zhu et al. 2022), which are specifically designed for document-level reasoning across multiple pages. To further enhance visual diversity, we also include SlideVQA (Tanaka et al. 2023), a dataset based on multi-slide presentations, and Multihiertt (Zhao et al. 2022), which focuses on complex hierarchical visual structures in multi-page charts. In addition, academic paper reading is a crucial task in the multi-page document domain. To improve the model’s capability in this area, we annotate a training subset, denoted as $\text{ArxivFullQA}_{\text{train}}$, from the DocMatrix dataset (Laurençon et al. 2024), following the same annotation details described in the next section.
@@
A key application of multi-page document understanding is the comprehension of scientific papers. However, large-scale multi-page benchmarks are still limited. To fill this gap, we curate a subset from the large-scale DocMatrix dataset (Laurençon et al. 2024), which contains scientific papers sourced from ArXiv. Unlike EviBench, the annotation details for ArxivFullQA differ in two key aspects. First, as shown in the supplementary material, the prompt used for annotation is specifically tailored to scientific papers. Second, in step 1, the input consists solely of LaTeX-formatted text extracted from DocMatrix, rather than image-based document representations, in order to improve the accuracy of the generated annotations. This enriched textual input guides the model in generating QA pairs across seven categories: factual, reasoning, comparison, summary, procedural, motivation, and result questions. In Step 2, the input switches from LaTeX text to the visual format of the full paper, and the question is taken from the QA pair annotated in Step 1. The data is retained only if the model’s answer matches the previously annotated answer. Following this process, we curate a new evaluation dataset, ArxivFullQA, comprising 8.6k high-quality multi-page QA samples.
@@
We initialize our training with the Qwen2.5-VL-Instruct model (Bai et al. 2025) for two main reasons. First, large-scale chain-of-thought training data for multi-page document understanding is extremely scarce, and collecting such annotations is highly resource-intensive. Second, this instruction-tuned model already exhibits a moderate level of reasoning ability, making it a practical alternative to the conventional “cold start” phase in GRPO. To further adapt the model to our task, we adopt a two-stage curriculum training strategy within the EviGRPO framework, as illustrated in Figure 2. In the first stage, we train the model for one epoch using only single-page data. This step activates its latent reasoning capability while aligning its outputs with the expected answer format. In the second stage, we continue training on multi-page data for another epoch to enhance the model’s ability to reason over longer contexts and across multiple pages.
@@
We evaluate our approach using both single-page and multi-page benchmarks. For single-page evaluation, we consider six widely used datasets: DocVQA (Mathew, Karatzas, and Jawahar 2021), InfographicVQA (Mathew et al. 2022), WikiTableQuestions (Pasupat and Liang 2015), TabFact (Chen et al. 2019), TextVQA (Singh et al. 2019), and VisualMRC (Tanaka, Nishida, and Yoshida 2021). For multi-page document understanding, we assess our method on six publicly available benchmarks: DUDE (Van Landeghem et al. 2023), MP-DocVQA (Tito, Karatzas, and Valveny 2023), TATDoc (Zhu et al. 2022), Multi-ChartQA (Zhu et al. 2024), MultiHiertt (Zhao et al. 2022), and SlideVQA (Tanaka et al. 2023), as well as our newly annotated benchmark, ArxivFullQA, to comprehensively evaluate its effectiveness in processing complex, multi-page inputs. Detailed dataset statistics are provided in the supplementary material.
@@ Appendix A Evaluation Benchmarks
To evaluate the effectiveness of our approach, we select six single-page and seven multi-page document understanding benchmarks. Notably, for datasets whose official test sets are not publicly accessible—TextVQA, DUDE, MP-DocVQA, and MultiHiertt—we follow prior works (Jia et al. 2024; Zhou et al. 2024) and report results on their validation sets.

### v3

@@ Introduction
Documents are ubiquitous in everyday life, encompassing scanned forms, tables, charts, and PDFs. Consequently, document understanding is a critical capability for multimodal large language models (MLLMs). While recent MLLMs have achieved strong performance on single-page document tasks (hu2024mplug1.5; zhou2024doge; liu2024textmonkey; wang2023towards; zhao2024tabpedia; feng2024docpedia; Xia_2024; dong2025spreadsheetllmencodingspreadsheetslarge), these tasks often cover a subset of real-world applications. In practice, many tasks involve multi-page documents—such as scientific papers, contracts, and reports—which require not only fine-grained visual understanding but also the ability to retrieve relevant pages and reason across dispersed content.
@@
Recent advances, including OpenAI’s o1 (jaech2024openai) and DeepSeek-R1 (guo2025deepseek), highlight the growing importance of advanced reasoning in LLMs. Parallel work has begun to explore reinforcement learning (RL) in MLLMs. For example, VisualRFT (liu2025visual) introduced verifiable reward functions for perception tasks and applied GRPO (shao2024deepseekmath) to enhance object detection and visual classification. Similarly, Vision-R1 (huang2025vision) integrated GRPO with a PTST training strategy to improve mathematical reasoning capabilities. However, these methods are restricted to single-image inputs and focus on either perception or symbolic reasoning. For multi-image tasks, methods like Video-R1 (feng2025video) and VideoChat-R1 (li2025videochat) have applied GRPO in the video domain, incorporating customized reward functions to capture spatiotemporal dependencies. Nevertheless, applying RL to multi-page document understanding—where effective evidence retrieval and multi-hop reasoning are both critical—remains largely unexplored.
@@
MLLMs have recently advanced document understanding by enabling unified modeling of textual and visual information, without relying on traditional OCR engines. Several recent works have explored MLLM-based document understanding. UReader (ye2023ureader) pioneered OCR-free, end-to-end instruction tuning across multiple tasks. The mPLUG-DocOwl series (ye2023mplug; hu2024mplug1.5; hu2024mplug2) improved structure awareness and computational efficiency through structural modeling and visual token compression. TextMonkey (liu2024textmonkey) introduced shifted window attention to preserve semantic continuity across image patches. DOGE (zhou2024doge) proposed a document-oriented grounding framework with high-quality training data, while Doc-CoB (mo2025doc) enabled dynamic region selection for focused, step-by-step reasoning. However, most methods are still limited to single-page documents. Our work targets more complex multi-page document tasks in real scenarios.
@@
RL has shown promise in enhancing the reasoning capabilities of LLMs. While early approaches relied on pretrained reward models, recent work such as DeepSeek-R1 (guo2025deepseek) demonstrates that simple, rule-based rewards can provide scalable and verifiable supervision. Building on this insight, researchers have extended RL to multimodal large language models (MLLMs). VisualRFT (liu2025visual) introduced visually grounded, verifiable reward functions for fine-grained classification and few-shot detection using GRPO. Vision-R1 (huang2025vision) further leveraged GRPO with a PTST strategy to improve mathematical reasoning on a curated multimodal dataset. In the video domain, Video-R1 (feng2025video) proposed T-GRPO to explicitly incorporate temporal cues during training, marking the first systematic application of GRPO to video-based MLLMs. VideoChat-R1 (li2025videochat) also applied GRPO to enhance spatiotemporal reasoning, achieving strong performance on temporal grounding and tracking tasks. While these studies focus on perception and video reasoning, RL has not yet been explored for multi-page document understanding. Our work aims to fill this gap.
@@
While the GRPO algorithm (shao2024deepseekmath) has demonstrated promise in enhancing reasoning capabilities, its direct application to multi-page document understanding remains suboptimal. To address this gap, we introduce Evidence Page-Guided GRPO (EviGRPO), a variant of GRPO tailored specifically for multi-page document tasks, as illustrated in Figure 2.
@@
As illustrated in Figure 3, we develop a rigorous two-stage data annotation pipeline. In the first stage, given an input and a task-specific prompt, we employ the Gemini 2.5 Flash model (comanici2025gemini) to generate a target output. The sample advances to the next stage only if its predicted answer is consistent with the ground truth. In the second stage, the same MLLM is prompted again with the annotated content in order to verify its accuracy. The annotation is retained only if the model’s output once again aligns with the ground truth under this controlled setting.
@@
As shown in Table 2, based on the proposed annotation pipeline, we presents the statistics of our annotated dataset. The dataset includes both single-page and multi-page document samples. Specifically, for single-page documents, we curate 13 widely-used datasets, including DocVQA (mathew2021docvqa), InfoGraphicsVQA (mathew2022infographicvqa), ChartQA (masry2022chartqa), DeepForm (svetlichnaya2020deepform), DVQA (kafle2018dvqa), FigureQA (kahou2017figureqa), KleisterCharity (stanislawek2021kleister), OCRVQA (mishra2019ocr), TabFact (chen2019tabfact), TextCaps (sidorov2020textcaps), TextVQA (singh2019towards), VisualMRC (tanaka2021visualmrc), and WikiTableQuestions (pasupat2015compositional). These datasets cover a diverse range of document types. For instance, DocVQA and DVQA contain scanned pages and visual charts, while DeepForm, TabFact, and WikiTableQuestions focus on structured tabular formats. This diversity enables the model to learn from various layouts and visual structures commonly encountered in real-world documents. For the multi-page document domain, we integrate three widely used datasets: DUDE (van2023document), MP-DocVQA (tito2023hierarchical), and TATDoc (zhu2022towards), which are specifically designed for document-level reasoning across multiple pages. To further enhance visual diversity, we also include SlideVQA (tanaka2023slidevqa), a dataset based on multi-slide presentations, and Multihiertt (zhao2022multihiertt), which focuses on complex hierarchical visual structures in multi-page charts. In addition, academic paper reading is a crucial task in the multi-page document domain. To improve the model’s capability in this area, we annotate a training subset, denoted as $\text{ArxivFullQA}_{\text{train}}$, from the DocMatrix dataset (laurenccon2024building), following the same annotation details described in the next section.
@@
A key application of multi-page document understanding is the comprehension of scientific papers. However, large-scale multi-page benchmarks are still limited. To fill this gap, we curate a subset from the large-scale DocMatrix dataset (laurenccon2024building), which contains scientific papers sourced from ArXiv. Unlike EviBench, the annotation details for ArxivFullQA differ in two key aspects. First, as shown in the supplementary material, the prompt used for annotation is specifically tailored to scientific papers. Second, in step 1, the input consists solely of LaTeX-formatted text extracted from DocMatrix, rather than image-based document representations, in order to improve the accuracy of the generated annotations. This enriched textual input guides the model in generating QA pairs across seven categories: factual, reasoning, comparison, summary, procedural, motivation, and result questions. In Step 2, the input switches from LaTeX text to the visual format of the full paper, and the question is taken from the QA pair annotated in Step 1. The data is retained only if the model’s answer matches the previously annotated answer. Following this process, we curate a new evaluation dataset, ArxivFullQA, comprising 8.6k high-quality multi-page QA samples.
@@
We initialize our training with the Qwen2.5-VL-Instruct model (bai2025qwen2.5vl) for two main reasons. First, large-scale chain-of-thought training data for multi-page document understanding is extremely scarce, and collecting such annotations is highly resource-intensive. Second, this instruction-tuned model already exhibits a moderate level of reasoning ability, making it a practical alternative to the conventional “cold start” phase in GRPO. To further adapt the model to our task, we adopt a two-stage curriculum training strategy within the EviGRPO framework, as illustrated in Figure 2. In the first stage, we train the model for one epoch using only single-page data. This step activates its latent reasoning capability while aligning its outputs with the expected answer format. In the second stage, we continue training on multi-page data for another epoch to enhance the model’s ability to reason over longer contexts and across multiple pages.
@@
We evaluate our approach using both single-page and multi-page benchmarks. For single-page evaluation, we consider six widely used datasets: DocVQA (mathew2021docvqa), InfographicVQA (mathew2022infographicvqa), WikiTableQuestions (pasupat2015compositional), TabFact (chen2019tabfact), TextVQA (singh2019towards), and VisualMRC (tanaka2021visualmrc). For multi-page document understanding, we assess our method on six publicly available benchmarks: DUDE (van2023document), MP-DocVQA (tito2023hierarchical), TATDoc (zhu2022towards), Multi-ChartQA (zhu2024multichartqa), MultiHiertt (zhao2022multihiertt), and SlideVQA (tanaka2023slidevqa), as well as our newly annotated benchmark, ArxivFullQA, to comprehensively evaluate its effectiveness in processing complex, multi-page inputs. Detailed dataset statistics are provided in the supplementary material.
@@ Conclusion
Category Dataset Images #Samples
Single DocVQA 1 5,188
InfographicVQA 1 3,288
TabFact 1 12,722
TextVQA 1 5,000
VisualMRC 1 6,729
WikiTableQuestions 1 4,343
Multi DUDE 1–20 11,402
MP-DocVQA 1-20 5019
TATDoc 1–2 1663
SlideVQA 20 2,136
MultiChartQA 2-3 2,000
Multihiert 3–7 7,830
ArxivFullQA 1–30 8,648
Table 7: Statistics of various benchmark datasets used for evaluation.
@@ Appendix A Evaluation Benchmarks
To evaluate the effectiveness of our approach, we select six single-page and seven multi-page document understanding benchmarks. Notably, for datasets whose official test sets are not publicly accessible—TextVQA, DUDE, MP-DocVQA, and MultiHiertt—we follow prior works (jia2024leopard; zhou2024doge) and report results on their validation sets.